# RBI-ThPhys-2025-26

### v2

@@ 3.3 Discreteness and Binning Effects
We will refer to this as the ‘effective wavenumber approximation’. This approximation will break down if the radial bins are too wide, or if the bispectrum varies strongly in the radial direction. One must exercise caution when attempting to apply these approximations (49). We tested their validity for our binning scheme by computing explicitly the sum over wavevectors and comparing to various choices for the wavenumbers $k_{i}^{\text{eff}}$ at which the bispectrum is evaluated. We also consider a comparison with numerical (Monte Carlo) integration over the wavevector bin, as done in e.g. (50; 9; 8). We then compute the SNR of the difference with respect to the explicit bin summation, using the diagonal sample variance as an error estimate, i.e.
$\text{SNR}^{L}_{\text{diff}}=\sum_{\text{bin }T}$ $(B_{\delta\delta E}^{L,\text{exact}}(T)-B_{\delta\delta E}^{L,\text{app.}}(T))C^{-1}_{LL}(T)$ (37)
@@
Here the ‘app.’ superscript refers to either the integral approximation or the effective wavenumber approximation and we do not consider cross-covariances between different multipoles. The result is shown in Figure 2. We can clearly see that using the effective wavenumber approximation (solid lines) is uncontrolled already at $k=0.10\,h/\text{Mpc}$. In contrast, the integral approximation (dashed lines) fares much better, with the SNR of the difference staying below unity for all relevant scales. We decide to perform exact binning of the theoretical predictions, as also done in (28). Here the situation may be more severe than for the scalar galaxy bispectrum, for which the multipoles exhibit weaker angular dependence, thus rendering the summand smoother and less susceptible to discretization effects [^8] It is also not clear whether the ordinary Legendre multipoles from Paper I would be less susceptible to discreteness effects; note that the summand in that case has a discontinuity along the $z$-axis where the phase factor $\exp 2\mathrm{i}\phi_{\mathbf{k}}$ is undefined..
@@ 3.4 Performance Metrics
where $L,L^{\prime}$ are multipole indices, $C_{LL^{\prime}}^{-1}$ is the precision matrix for a single realization and $\hat{\bar{B}}_{L}$ is the sample mean over all realizations as in Eq. (32). We suppress the tracers $XYZ$ and the triangle bins for clarity. Thus, we carry out our fiducial analysis at $V=1\,(\text{Gpc}/h)^{3}$ as in (28). We will explore several different sets of multipoles to be included in the likelihood in Section 4.
@@
Assuming that (i) the best-fit model $B_{L,*}^{\text{th}}$ found from minimizing $\bar{\chi}^{2}$ is the true mean of the distribution of $\hat{B}$ and (ii) the precision matrix $C_{LL^{\prime}}^{-1}$ equals the true precision matrix, the variable $\bar{\chi}^{2}_{*}=\bar{\chi}^{2}(B_{L,*}^{\text{th}})$ follows a $\chi^{2}$ distribution, which is approximately normal [^10] From now on, we will omit the asterisk on chi-squared statistics to avoid clutter. When referring to the distribution of e.g. Eq. (43), it is understood that it is evaluated at the best-fit value. For a more complete treatment of how this can impact e.g. the number of degrees of freedom, see (53).. When the covariance (and hence the precision matrix) is known exactly and one ignores the number of free parameters, which is much less than $N_{L}N_{T}$, one recovers the standard result
@@
Here $\mathcal{N}(\mu,\sigma^{2})$ denotes a normal distribution. Thus, we can assess the goodness of fit by checking whether the left-hand side is sufficiently close to unity for the best-fit model. If we assume the Gaussian covariance to be the true one, we can apply Eq. (45) without further issues. However, here an additional subtletly arises when the covariance is estimated from the same data, as with Eq. (32). In that case, the mean $\hat{\mu}_{\text{red}}$ and variance $\hat{\sigma}_{\text{red}}^{2}$ of $\hat{\bar{\chi}}^{2}/(N_{T}N_{L})$ (where the hat indicates that Eq. (32) is used for the covariance matrix) are not equal to the values from Eq. (45), even after applying the Hartlap correction from Eq. (33). Moreover, $\hat{\sigma}_{\mathrm{red}}^{2}$ does not approach $\sigma_{\mathrm{red}}^{2}$ even when $N_{r}\to\infty$. We elaborate on this issue in App. B, where we also derive the correct expressions for the mean and standard deviation of $\hat{\bar{\chi}}^{2}$ in this specific case. The chi-squared test can then still be used to assess the goodness of fit [^11] Strictly speaking, there are additional complications when using a noisy estimate for the precision matrix. In particular, the noise in the covariance matrix needs to be propagated into the likelihood inference as well (54). We ignore these issues here..
@@
We will compute all the corresponding quantities defined above in both the case where the Gaussian covariance is used, and the case where Eq. (32) is used, taking into account the above discussion. In order to be roughly consistent with (28), the FoB is normalized by its 68 percent confidence level. In (28), a threshold of $\Delta_{\text{crit}}=1.5$ was used to define the breakdown of the model (see also (52; 55)). In this work, given the uncertainties we face when computing the $\chi^{2}$ statistic [^13] This is due to the uncertainty in the bispectrum covariance; the large number of degrees of freedom renders the p-value for the $\chi_{\text{red}}^{2}$ statistic sensitive to even $\sim 5\%$ differences in the covariance. we opt to be somewhat more lenient and use $\Delta_{\text{crit}}=2$.
@@ 4 Results
We used the diagonal sample variance to compute the SNR, but found no appreciable difference when using the Gaussian covariance instead. The result is shown in Figure 3, where we clearly see in the top panel that all parity-even multipoles are detected at $k_{\text{max}}=0.11\,h/\text{Mpc}$ (which will turn out to be the wavenumber at which our model breaks down), with the monopole $\tilde{B}^{00}_{\delta\delta E}$ having the highest SNR. All parity-odd multipoles are also detected at this wavenumber, except for $B_{BBB}^{22}$, which we indeed expect to be zero at tree-level in the EFT of IA (see the discussion around Eq. (23)). This also serves as an important consistency check for our parity-odd estimators; the inclusion of the cross-product in Eq. (24) allows us to extract the 3D bispectrum signal after averaging over all triangle orientations.
Figure 5: Comparison of the posteriors from bispectra with at most one (green), two (grey) or three (red) shape fields at $k_{\text{max}}=0.11\,h/\text{Mpc}$. Also show is the (Gaussian) power spectrum posterior from (28), cf. Eq. (39) and (40) (where a prior on the sign of $b_{2,2}^{\text{g}}$ and $b_{2,3}^{\text{g}}$ had to be applied). Observe that the grey and red posteriors are almost identical.
@@
Figure 6: Posteriors for the first and second order IA bias parameters at $k_{\text{max}}=0.11\,h/\text{Mpc}$ when analyzing individual multipoles of $B_{\delta\delta E}$ and when combining them. It is apparent that each multipole individually suffers from severe degeneracies, but combining them (blue contours) yields very good constraints on their amplitudes.
@@
Figure 8: Measured bispectrum signal for the four multipoles with one tensor field considered in this work. We plot the product $k_{1}k_{2}k_{3}\times B$ in units of $(\text{Mpc}/h)^{3}$. Each row contains a specific multipole, while each column contains a specific type of triangle configuration. In the left column, we plot ‘obtuse’ configurations where the sum of the two shortest sides exceeds the longest by roughly $0.02\,h/\text{Mpc}$. Note that these plots are therefore symmetric about the ‘middle bin’ for the even multipoles (first three rows); that is, the same data is plotted twice, but this is not true for $B_{\delta\delta B}^{00}$. In the leftmost and rightmost columns, we choose two specific values for the remaining ‘fixed’ side $k_{1}$ (pink and blue). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.
where the respective lines indicate bispectra with one, two or three tensor fields (cf. Eq. (22)). Note that for the second line, the $(\ell_{1},\ell_{2})=(0,2)$ multipole is in fact the lowest order non-vanishing one, and a similar remark applies to the third line. We will discuss the impact of adding each line to the likelihood extensively in the three subsections below, i.e. Section 4.1 deals with bispectra with one shape field, while Section 4.2 deals with the five bispectra with at most two shape fields and Section 4.3 discusses the likelihood with all seven bispectra above. Our main results are summarized in Figures 4 and 5. In Figure 4, we see that the bispectrum is nominally well described by the EFT of IA up to $k_{\text{max}}=0.11\,h/\text{Mpc}$ (vertical dashed line), with some variations depending on the specific setup. For $k_{\text{max}}=0.11\,h/\text{Mpc}$, the posteriors on the three parameters included in the FoB metric from Eq. (41) are plotted in Figure 5, along with the power spectrum result from (28), indicating good consistency between two- and three-point statistics. The bispectrum posteriors are still broader than those from the power spectrum, but given that the power spectrum analysis used a much larger value of $k_{\text{max}}=0.28\,h/\text{Mpc}$ this is not surprising.
@@ 4.1 One shape field
Figure 9: Same as Figure 8 but now for multipoles with two shape fields. The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.
@@ 4.2 One or two shape fields
Figure 11: Same as Figure 8 but now for multipoles with three shape fields. Observe again that the odd bispectra vanish in the equilateral configuration (Figure 6 of Paper I) because of symmetries (and the fact that two of the fields are the same). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51); we stress that these are not fits to the data.
@@ 4.3 One, two or three shape fields
Previously, the power spectrum-only analysis of (28) was complicated to some extent by the quadratic dependence of the model predictions on the second order bias parameters. In that work, parameter posteriors of the second order bias parameters $b_{2,1}^{\text{g}}$ and $b_{2,3}^{\text{g}}$ were found to exhibit bimodalities when analyzing only the one-loop power spectrum of IA. To mitigate this issue, (28) assigned a physically motivated prior on the signs of these parameters as expected from a local Lagrangian bias assumption (58; 59). We showed that this assumption is correct and that the degeneracies present in the one-loop power spectrum analysis can be broken through the addition of the bispectrum. Indeed, in the tree-level bispectrum, all second order bias parameters appear linearly in the model (in fact, in the case of e.g. $B_{\delta\delta E}$ there is only one biased tracer, so all parameters appear linearly in the model). Thus, the degeneracies present in the power spectrum likelihood are guaranteed to be broken. Our conclusive determination of the second order bias parameters has some immediate interesting phenomenological applications. For example, in terms of the more familiar second order basis $\{b_{\delta K},b_{K^{2}},b_{t}\}$ (see e.g. (28) for definitions of these operators and the corresponding basis transformation) we obtain the parameter constraints
@@ 5 Conclusions and Discussion
Note that simpler models for the scale dependence of intrinsic alignment bispectra fall short here; for example according to NLA-inspired models (19; 36) there is no B-mode and we would have $B_{\delta\delta B}=0$. However, we have clearly demonstrated that this bispectrum is nonzero. In addition, our work indicates that bispectra with more than one shape field are strongly dependent on the noise amplitude $\alpha_{\text{gg}}^{1\delta}$. This contribution was missing in the earlier second-order (i.e. ‘TATT’) alignment models of (35; 60). We emphasize that the importance of this contribution will only be exacerbated as we consider more realistic samples with lower number densities and lower alignment amplitudes. To this end, it is reassuring that several different bispectra can in principle be combined in order to jointly constrain noise amplitudes. This is also important when attempting to use the bispectrum of IA for mitigation purposes in cosmic shear surveys. Generally speaking, other models for IA which are inspired by the galaxy bias expansion are also capable of providing physically motivated templates for the IA bispectrum that can be used in conjunction with two-point statistics. This may allow for more efficient nuisance parameter choices (59). Alternatively, using the Lagrangian framework, (61; 62) showed that a perturbative description of intrinsic alignments can be used to place conservative constraints on the amplitude of matter fluctuations $S_{8}\equiv\sigma_{8}(\Omega_{m}/0.3)^{0.5}$.We anticipate that the bispectrum of IA can be a useful tool in directly constraining the intrinsic alignment of galaxy samples when spectroscopic redshifts are available (26; 27).
@@
Our results motivate a more in-depth study of joint constraints wrought from the power spectrum and bispectrum of IA for a more realistic halo or galaxy sample. This can be done by downsampling halos from an N-body simulation and randomizing their orientations to mimic misalignment of the LRG with respect to its host (63; 64), or by considering a galaxy population from a large-scale hydrodynamical simulation such as (65). We will explore these options in future work.

### v3

@@ 3.3 Discreteness and Binning Effects
We will refer to this as the ‘effective wavenumber approximation’. This approximation will break down if the radial bins are too wide, or if the bispectrum varies strongly in the radial direction. One must exercise caution when attempting to apply these approximations (55). We tested their validity for our binning scheme by computing explicitly the sum over wavevectors and comparing to various choices for the wavenumbers $k_{i}^{\text{eff}}$ at which the bispectrum is evaluated. We also consider a comparison with numerical (Monte Carlo) integration over the wavevector bin, as done in e.g. (56; 9; 8). We then compute the SNR of the difference with respect to the explicit bin summation, using the diagonal sample variance as an error estimate, i.e.
$(\text{SNR}^{L}_{\text{diff}})^{2}=\sum_{\text{bin }T}$ $(B_{\delta\delta E}^{L,\text{exact}}(T)-B_{\delta\delta E}^{L,\text{app.}}(T))C^{-1}_{LL}(T)$ (37)
@@
Here the ‘app.’ superscript refers to either the integral approximation or the effective wavenumber approximation and we do not consider cross-covariances between different multipoles. The result is shown in Figure 2. We would like for this SNR to remain small, i.e. below unity. We can clearly see that using the effective wavenumber approximation (dashed lines) is uncontrolled already at $k=0.10\,h/\text{Mpc}$. In contrast, the integral approximation (solid lines) fares much better, with the SNR of the difference staying below unity for all relevant scales. We decide to perform exact binning of the theoretical predictions, as also done in (31). Here the situation may be more severe than for the scalar galaxy bispectrum, for which the multipoles exhibit weaker angular dependence, thus rendering the summand smoother and less susceptible to discreteness effects [^8] It is also not clear whether the ordinary Legendre multipoles from Paper I would be less susceptible to discreteness effects; note that the summand in that case has a discontinuity along the $z$-axis where the phase factor $\exp 2\mathrm{i}\phi_{\mathbf{k}}$ is undefined..
@@ 3.4 Performance Metrics
where $L,L^{\prime}$ are multipole indices, $C_{LL^{\prime}}^{-1}$ is the precision matrix for a single realization and $\hat{\bar{B}}_{L}$ is the sample mean over all realizations as in Eq. (32). We suppress the tracers $XYZ$ and the triangle bins for clarity. Thus, we carry out our fiducial analysis at $V=1\,(\text{Gpc}/h)^{3}$ as in (31). By using the mean of the data vectors over all realizations, the statistical noise becomes suppressed with respect to the covariance used. This allows us to more cleanly identify modeling systematics at a given scale cut. We will explore several different sets of multipoles to be included in the likelihood in Section 4.
@@
Assuming that (i) the best-fit model $B_{L,*}^{\text{th}}$ found from minimizing $\bar{\chi}^{2}$ is the true mean of the distribution of $\hat{B}$ and (ii) the precision matrix $C_{LL^{\prime}}^{-1}$ equals the true precision matrix, the variable $\bar{\chi}^{2}_{*}=\bar{\chi}^{2}(B_{L,*}^{\text{th}})$ follows a $\chi^{2}$ distribution, which is approximately normal [^10] From now on, we will omit the asterisk on chi-squared statistics to avoid clutter. When referring to the distribution of e.g. Eq. (43), it is understood that it is evaluated at the best-fit value. For a more complete treatment of how this can impact e.g. the number of degrees of freedom, see (59).. When the covariance (and hence the precision matrix) is known exactly and one ignores the number of free parameters, which is much less than $N_{L}N_{T}$ (where $N_{T}$ is the number of triangle bins), one recovers the standard result
@@
Here $\mathcal{N}(\mu,\sigma^{2})$ denotes a normal distribution. Thus, we can assess the goodness of fit by checking whether the left-hand side is sufficiently close to unity for the best-fit model. If we assume the Gaussian covariance to be the true one, we can apply Eq. (45) without further issues. However, here an additional subtlety arises when the covariance is estimated from the same data, as with Eq. (32). In that case, the mean $\hat{\mu}_{\text{red}}$ and variance $\hat{\sigma}_{\text{red}}^{2}$ of $\hat{\bar{\chi}}^{2}/(N_{T}N_{L})$ (where the hat indicates that Eq. (32) is used for the covariance matrix) are not equal to the values from Eq. (45), even after applying the Hartlap correction from Eq. (33). Moreover, $\hat{\sigma}_{\mathrm{red}}^{2}$ does not approach $\sigma_{\mathrm{red}}^{2}$ even when $N_{r}\to\infty$. We elaborate on this issue in App. B, where we also derive the correct expressions for the mean and standard deviation of $\hat{\bar{\chi}}^{2}$ in this specific case. The chi-squared test can then still be used to assess the goodness of fit [^11] Strictly speaking, there are additional complications when using a noisy estimate for the precision matrix. In particular, the noise in the covariance matrix needs to be propagated into the likelihood inference as well (60). We ignore these issues here..
@@
We will compute all the corresponding quantities defined above in both the case where the Gaussian covariance is used, and the case where Eq. (32) is used, taking into account the above discussion. In order to be roughly consistent with (31), the FoB is normalized by its 68 percent confidence level. In (31), a threshold of $\Delta_{\text{crit}}=1.5$ was used to define the breakdown of the model (see also (58; 61)). In this work, given the uncertainties we face when computing the $\chi^{2}$ statistic [^13] This is due to the uncertainty in the bispectrum covariance; the large number of degrees of freedom renders the p-value for the $\chi_{\text{red}}^{2}$ statistic sensitive to even $\sim 5\%$ differences in the covariance. we opt to be somewhat more lenient and use $\Delta_{\text{crit}}=2$.
@@ 4 Results
We used the diagonal sample variance to compute the SNR, but found no appreciable difference when using the Gaussian covariance instead. The result is shown in Figure 3, where we clearly see in the top panel that all parity-even multipoles are detected at $k_{\text{max}}=0.11\,h/\text{Mpc}$ (which will turn out to be the wavenumber at which our model breaks down), with the monopole $\tilde{B}^{00}_{\delta\delta E}$ having the highest SNR. All parity-odd multipoles are also detected at this wavenumber, except for $B_{BBB}^{22}$, which we indeed expect to be zero at tree-level in the EFT of IA (see the discussion around Eq. (23)). This also serves as an important consistency check for our parity-odd estimators; the inclusion of the cross-product in Eq. (24) allows us to extract the 3D bispectrum signal after averaging over all triangle orientations. It is important to note that these SNR values are obtained without performing any kind of ‘noise’ subtraction from the measured bispectrum, that is, the stochastic contributions proportional to $\alpha_{\text{gg}}^{1\delta}$ in $B^{\delta\text{gg}}$ and $B^{\text{ggg}}$ are not removed. These contributions constitute a large part of the measured bispectra, and we will come back to this issue in Sections 4.2 and 4.3. Importantly (and in line with general expectation), the SNR does not directly translate into the quality of parameter constraints.
Figure 5: Comparison of the posteriors from bispectra with at most one (green), two (grey) or three (red) shape fields at $k_{\text{max}}=0.11\,h/\text{Mpc}$. Also show is the (Gaussian) power spectrum posterior from (31), cf. Eq. (39) and (40) (where a prior on the sign of $b_{2,2}^{\text{g}}$ and $b_{2,3}^{\text{g}}$ had to be applied). Observe that the grey and red posteriors are almost identical. Here we used the Hartlap-corrected diagonal sample variance from Eq. (32)
@@
Figure 6: Posteriors for the first and second order IA bias parameters at $k_{\text{max}}=0.11\,h/\text{Mpc}$ when analyzing individual multipoles of $B_{\delta\delta E}$ and when combining them. It is apparent that each multipole individually suffers from severe degeneracies, but combining them (blue contours) yields very good constraints on their amplitudes. Here we used the Hartlap-corrected diagonal sample variance from Eq. (32).
@@
Figure 8: Measured bispectrum signal for the four multipoles with one tensor field considered in this work. We plot the product $k_{1}k_{2}k_{3}\times B$ in units of $(\text{Mpc}/h)^{3}$. Each row contains a specific multipole, while each column contains a specific type of triangle configuration. In the left column, we plot ‘obtuse’ configurations where the sum of the two shortest sides exceeds the longest by roughly $0.02\,h/\text{Mpc}$. Note that these plots are therefore symmetric about the ‘middle bin’ for the even multipoles (first three rows); that is, the same data is plotted twice, but this is not true for $B_{\delta\delta B}^{00}$. In the leftmost and rightmost columns, we choose two specific values for the remaining ‘fixed’ side $k_{1}$ (pink and blue). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51). The bispectra have been multiplied by $k_{123}=k_{1}k_{2}k_{3}$ for visualization purposes.
where the respective lines indicate bispectra with one, two or three tensor fields (cf. Eq. (22)). Note that for the second line, the $(\ell_{1},\ell_{2})=(0,2)$ multipole is in fact the lowest order non-vanishing one, and a similar remark applies to the third line. We will discuss the impact of adding each line to the likelihood extensively in the three subsections below, i.e. Section 4.1 deals with bispectra with one shape field, while Section 4.2 deals with the five bispectra with at most two shape fields and Section 4.3 discusses the likelihood with all seven bispectra above. All fits are shown both for the use of the (Hartlap-corrected) diagonal sample variance from Eq. (32) as well as the theoretical Gaussian covariance. Our main results are summarized in Figures 4 and 5. In Figure 4, we see that the bispectrum is nominally well described by the EFT of IA up to $k_{\text{max}}=0.11\,h/\text{Mpc}$ (vertical dashed line), with some variations depending on the specific setup. For $k_{\text{max}}=0.11\,h/\text{Mpc}$, the posteriors on the three parameters included in the FoB metric from Eq. (41) are plotted in Figure 5, along with the power spectrum result from (31), indicating good consistency between two- and three-point statistics. The bispectrum posteriors are still broader than those from the power spectrum, but given that the power spectrum analysis used a much larger value of $k_{\text{max}}=0.28\,h/\text{Mpc}$ this is not surprising.
@@ 4.1 One shape field
Figure 9: Same as Figure 8 but now for multipoles with two shape fields. The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51). The bispectra have been multiplied by $k_{123}=k_{1}k_{2}k_{3}$ for visualization purposes.
@@ 4.2 One or two shape fields
Figure 11: Same as Figure 8 but now for multipoles with three shape fields. Observe again that the odd bispectra vanish in the equilateral configuration (Figure 6 of Paper I) because of symmetries (and the fact that two of the fields are the same). The errors have been rescaled to match a volume of $V=7.5\,(\text{Gpc}/h)^{3}$. The dashed lines show the theory prediction using the values from Eq. (51). The bispectra have been multiplied by $k_{123}=k_{1}k_{2}k_{3}$ for visualization purposes.
@@ 4.3 One, two or three shape fields
Previously, the power spectrum-only analysis of (31) was complicated to some extent by the quadratic dependence of the model predictions on the second order bias parameters. In that work, parameter posteriors of the second order bias parameters $b_{2,1}^{\text{g}}$ and $b_{2,3}^{\text{g}}$ were found to exhibit bimodalities when analyzing only the one-loop power spectrum of IA. To mitigate this issue, (31) assigned a physically motivated prior on the signs of these parameters as expected from a local Lagrangian bias assumption (64; 65). We showed that this assumption is correct and that the degeneracies present in the one-loop power spectrum analysis can be broken through the addition of the bispectrum. Indeed, in the tree-level bispectrum, all second order bias parameters appear linearly in the model (in fact, in the case of e.g. $B_{\delta\delta E}$ there is only one biased tracer, so all parameters appear linearly in the model). Thus, the degeneracies present in the power spectrum likelihood are guaranteed to be broken. Our conclusive determination of the second order bias parameters has some immediate interesting phenomenological applications. For example, in terms of the more familiar second order basis $\{b_{\delta K},b_{K^{2}},b_{t}\}$ (see e.g. (31) for definitions of these operators and the corresponding basis transformation) we obtain the parameter constraints
@@ 5 Conclusions and Discussion
Note that simpler models for the scale dependence of intrinsic alignment bispectra fall short here; for example according to NLA-inspired models (21; 22) there is no B-mode and we would have $B_{\delta\delta B}=0$. However, we have clearly demonstrated that this bispectrum is nonzero. In addition, our work indicates that bispectra with more than one shape field are strongly dependent on the noise amplitude $\alpha_{\text{gg}}^{1\delta}$. This contribution was missing in the earlier second-order (i.e. ‘TATT’) alignment models of (39; 66). We emphasize that the importance of this contribution will only be exacerbated as we consider more realistic samples with lower number densities and lower alignment amplitudes. To this end, it is reassuring that several different bispectra can in principle be combined in order to jointly constrain noise amplitudes. This is also important when attempting to use the bispectrum of IA for mitigation purposes in cosmic shear surveys. Generally speaking, other models for IA which are inspired by the galaxy bias expansion are also capable of providing physically motivated templates for the IA bispectrum that can be used in conjunction with two-point statistics. This may allow for more efficient nuisance parameter choices (65). Alternatively, using the Lagrangian framework, (32; 67) showed that a perturbative description of intrinsic alignments can be used to place conservative constraints on the amplitude of matter fluctuations $S_{8}\equiv\sigma_{8}(\Omega_{m}/0.3)^{0.5}$.We anticipate that the bispectrum of IA can be a useful tool in directly constraining the intrinsic alignment of galaxy samples when spectroscopic redshifts are available (33; 34).
@@
Our results motivate a more in-depth study of joint constraints wrought from the power spectrum and bispectrum of IA for a more realistic halo or galaxy sample. This can be done by downsampling halos from an N-body simulation and randomizing their orientations to mimic misalignment of the LRG with respect to its host (68; 69), or by considering a galaxy population from a large-scale hydrodynamical simulation such as (70). We will explore these options in future work.